# A comprehensive framework for multi-fidelity surrogate modeling with noisy data: a gray-box perspective

### v1

@@ Abstract
Computer simulations (a.k.a. white-box models) are more indispensable than ever to model intricate engineering systems. However, computational models alone often fail to fully capture the complexities of reality. When physical experiments are accessible though, it is of interest to enhance the incomplete information offered by computational models. Gray-box modeling is concerned with the problem of merging information from data-driven (a.k.a. black-box) models and white-box (i.e., physics-based) models. In this paper, we propose to perform this task by using multi-fidelity surrogate models (MFSMs). A MFSM integrates information from models with varying computational fidelities into a new surrogate model. The multi-fidelity surrogate modeling framework we propose handles noise-contaminated data and is able to estimate the underlying noise-free high-fidelity function. Our methodology emphasizes on delivering precise estimates of the uncertainty in its predictions in the form of confidence and prediction intervals, by quantitatively incorporating the different types of uncertainty that affect the problem, arising from measurement noise and from lack of knowledge due to the limited experimental design budget on both the high- and low-fidelity models. Applied to gray-box modeling, our MFSM framework treats noisy experimental data as the high-fidelity and the white-box computational models as their low-fidelity counterparts. The effectiveness of our methodology is showcased through synthetic examples and a wind turbine application.
@@ 1 Introduction
Predicting the behavior of complex systems and quantifying the corresponding uncertainty is a ubiquitous challenge in engineering and applied sciences. Two essential tools that engineers and scientists use to deal with this challenge are experiments and predictive models. Generally, experiments can be classified into two broad categories: physical and computer experiments. The latter, also known as computer simulations, are mathematical or computational models of a system which use pre-existing knowledge on the underlying physics of the system to predict its behaviour. Predictive models of this type can be referred to as white-box models (Rogers et al., 2017).
Although white-box models are interpretable, they sometimes fail to capture the entirety of the system they represent due to simplifying assumptions and approximations often needed to make them computationally tractable. A second class of predictive models are the so-called black-box models. These consist in data-driven models that act as global approximators of the response of a system, based on an available set of input-output observations. While black-box models can provide flexibility, they may not honor the underlying physics as their white-box counterparts do (Rogers et al., 2017).
Often, the response of a system can be predicted by one or more white-box computational models, while additional data can be obtained through physical experiments. In this context, it is natural to look for an approach that combines the white- and black-box modeling paradigms to benefit from the strengths of both. This combined approach is known as gray-box modeling (Tulleken, 1993). Traditionally, physical experiments have been used to improve computational models through model calibration, where parameters of a model are inferred by fitting the model predictions to available experimental data (Kennedy and O'Hagan, 2001; Higdon et al., 2004). Gray-box models take a more comprehensive approach by acknowledging that the computational model may not fully capture the system’s complexity and thus, incorporating both knowledge-driven and data-driven elements. As an example, hybrid simulation, which combines physical and numerical substructures to create a hybrid model (Schellenberg et al., 2009; Abbiati et al., 2021), follows a gray-box modeling paradigm. Recently, machine learning approaches using physics-informed neural networks (Raissi et al., 2019) have been employed to perform gray-box modeling, e.g. in the work of Yan et al. (2022). Finally, gray-box models have been historically employed in the field of control theory and system identification, where experimental data are used to estimate the parameters and structure of a mathematical model that represents the underlying dynamical system (Ljung, 1998).
In many applications which require a large number of model evaluations, such as uncertainty quantification and optimization, an obstacle frequently encountered is the high computational cost associated with white-box simulations. For this reason, computational models are often replaced by surrogate models (SMs), also known as metamodels, or emulators. A SM acts as an inexpensive-to-evaluate approximator of an original model, and is constructed using a limited set of model evaluations, called the experimental design (ED), also known as training set in machine learning. Among the most widely used SMs for deterministic simulators are polynomial chaos expansions (PCE) (Xiu and Karniadakis, 2002; Blatman and Sudret, 2011), Gaussian processes (GPs) (Rasmussen and Williams, 2006), and support vector regression (Drucker et al., 1996).
@@
The choice of surrogate model is an integral part of the design and construction of a MFSM. Numerous MFSM techniques are based on Gaussian process modeling, following the autoregressive fusing scheme proposed by Kennedy and O’Hagan (2000). Such works include Forrester et al. (2007); Kuya et al. (2011), and Le Gratiet and Garnier (2014), among others, with the latter reformulating the approach from Kennedy and O’Hagan (2000) to have a recursive form, allowing for a reduced computational complexity. Polynomial chaos expansion is another SM that has gained popularity in the past two decades for the purpose of MF surrogate modeling (Ng and Eldred, 2012; Palar et al., 2016). The approaches mentioned so far use linear information fusion, which entail assuming that a higher-fidelity response can be expressed as a linear combination of a lower-fidelity model and a discrepancy function. Recently, multi-fidelity modeling approaches have been proposed in the machine learning community, for example the deep GP-based framework of Cutajar et al. (2018), the Bayesian neural network approaches from Meng et al. (2021); Kerleguer et al. (2024), and the generative adversarial network-based methodology proposed by Zhang et al. (2022). These approaches can capture the nonlinear relations between the different levels of fidelity. According to a comparison among different linear and nonlinear GPs-based MF techniques performed by Brevault et al. (2020), when the high- and low-fidelity models are weakly correlated, nonlinear techniques can outperform linear and less complex techniques, with the caveat of requiring a larger quantity of HF data.
In this paper, we propose to perform gray-box modeling using multi-fidelity surrogate models. For this purpose, we assume that the white-box computational models can capture the general behavior of the physics of the system, while the experiments can capture its entirety, but only on a very limited set of conditions, due to their associated costs. In particular, we consider the experimental data as noise-contaminated realizations of black-box HF models, whereas the available white-box computational models are considered as their LF counterparts. Because white-box computational models can in principle be expensive to evaluate, we also approximate them with surrogate models, constructed on a dataset of low-fidelity model evaluations. White-box models are approached within a nonintrusive context, eliminating the requirement for prior knowledge of the underlying equations.
@@
Most of the existing literature on MF surrogate modeling focuses on deterministic and noise-free high- and low-fidelity models. However, all real world measurement devices have a limited precision and resolution, and therefore, data resulting from measurements are generally contaminated by irriducible noise. This noise can be considered as a source of aleatory uncertainty, as it is inherent in the measurement process. Moreover, the available training data for the construction of all the elements of a MFSM is in principle relatively small due to computational budget constraints. Thus, we consider all MFSM predictions as affected by epistemic uncertainty. Recent studies that consider the presence of noise in a MF setting include the work of Raissi et al. (2017), who use GP regression to infer the solutions of differential equations when noisy data of different fidelities are available. Furthermore, Zhang et al. (2018) demonstrate that their linear regression-based MF surrogate modeling technique is robust to noise in the HF data, and is also able to estimate the noise level when enough HF data is available. Finally, Ficini et al. (2021) assess the robustness of a GP regression MFSM on problems affected by noisy objective function evaluations.
@@
To the authors’ best knowledge, however, no work has yet introduced a comprehensive approach to multi-fidelity surrogate modeling that considers the presence of both noise and/or epistemic uncertainty in the high- and/or low-fidelity data. This approach should also provide a way to quantify the accuracy of the MF model predictions with respect to a) the underlying noise-free HF model, and b) the noisy HF observations. The uncertainty about the mean prediction of a regression model is typically expressed via confidence intervals (CIs). On the other hand, the uncertainty on the prediction of an unseen noise-contaminated observation is shown via prediction intervals (PIs) (Kutner et al., 2005). Generally, PIs are wider than the corresponding CIs. Many of the existing MF surrogate modeling methods do not provide uncertainty estimations about the MF model predictions. In the cases when CIs are provided, their construction is linked to the particular methodology employed in constructing the MFSM, e.g., GP-based CIs (Raissi et al., 2017; Perdikaris et al., 2017). This can potentially make the frameworks less flexible when different MFSM architectures or SMs need to be explored. Moreover, no work has considered the distinction between confidence and prediction intervals in a MF setting; typically, the latter are disregarded.
@@
The goal of the present paper is to introduce a novel general framework for multi-fidelity surrogate modeling that is able to deal with noisy data and epistemic uncertainty due to limited training information. We assume that the noisy HF data originate from deterministic models contaminated by unbiased stochastic noise. Therefore our MF framework aims at effectively emulating the underlying deterministic models, or in other words at denoising the black-box component. An essential and original feature of our methodology is its ability to provide estimates of the different kinds of uncertainty in its predictions in the form of both confidence and prediction intervals. Moreover, our framework can be applied in the field of gray-box modeling, where experimental data and computational models can be fused into a gray-box multi-fidelity surrogate model.
@@
This paper is organized as follows: In Section 2, after recalling the relevant theory for MFSMs, we present our framework for MF gray-box modeling, including the related confidence and prediction intervals, as well as the implementation details for each component of the framework. Subsequently, the proposed method is assessed in Section 3, where it is applied on two synthetic examples of increasing complexity and on a real-world application. Finally, in Section 4 we discuss concluding remarks and present prospects for future research.
@@ 2 Methods
In this section, we first formally state the multi-fidelity surrogate modeling problem, while establishing the notation we are adopting throughout the paper. Then, we describe our proposed methodology to construct a gray-box multi-fidelity surrogate model. Subsequently, we propose to express the uncertainty related to the MFSM predictions using confidence and prediction intervals, and we discuss the interpretation of these intervals in the MF setting. Lastly, we outline our proposed implementation for constructing a MF gray-box, as well as for estimating its confidence and prediction intervals.
@@ 2.1 Multi-fidelity surrogate modeling
The class of surrogates for the LF model and the discrepancy function can be chosen among a wide range of possibilities, including, among others, PCE, GP regression, or neural networks. However, the framework described in this section is independent of the particular choice of surrogate modeling methods, and hence, we will refrain from specifying the choice of SM for now.
@@
2.2 Gray-box modeling
Gray-box modeling consists in combining elements of both data-driven methods (black boxes), and physics-based computational models (white boxes) (Rogers et al., 2017). Gray-box models leverage existing physical knowledge about a system through their white-box elements, but due to their black-box components, they are also able to capture complex relationships in experimental data that might not be explicitly captured by white-box models. By combining data-driven methods with physical insights, they can compensate for gaps in the understanding of a system’s behavior. Hence, gray-box models can provide a more accurate and comprehensive representation of a system than either black- or white-box models alone.
Let us consider a set of observations $(\boldsymbol{{\mathcal{X}}}_{\text{H}},{\mathcal{Y}}_{\text{H}})$ obtained through an experimental campaign, each observation of which can be expressed as in Equation 3. The unknown underlying model $\psi_{\text{H}}$ can be considered as an expensive-to-evaluate black box. Moreover, let us assume that the system response can also be predicted by a white-box computational model, such as a system of equations or a complex finite-elements simulator, denoted as ${\mathcal{M}}_{\text{L}}(\boldsymbol{x})$.
Then, we can assume that experiments can accurately capture the behavior of the system under investigation, but the information they provide is incomplete, due to their scarcity. On the other hand, the available computational model can complement this information by providing a larger amount of data, but at the cost of lower accuracy, due either to inherent model simplifications (e.g. ignoring some physics), or to numerical limitations (e.g. discretization). Hence, it is possible to create a gray-box model using multi-fidelity surrogate modeling, wherein the experimental data is regarded as noise-affected realizations of a high-fidelity black-box model, and the white-box computational model is treated as its low-fidelity equivalent. Following the rationale described in the previous section, the computational model is replaced by a surrogate model ^$\widehat{\mathcal{M}}_{\text{L}}(\boldsymbol{x})$, constructed with a set of model evaluations $(\boldsymbol{{\mathcal{X}}}_{\text{L}},{\mathcal{M}}_{\text{L}}(\boldsymbol{{\mathcal{X}}}_{\text{L}}))$.
Adopting the MF surrogate modeling information fusion scheme introduced previously, a multi-fidelity gray-box model predictor aiming to emulate the underlying noise-free HF function output $\psi_{\text{H}}(\boldsymbol{x})$ can be expressed as:
@@ 2.3 Confidence and prediction intervals
Confidence and prediction intervals are powerful means of conveying the uncertainty present in the predictions of a model, thereby significantly elevating the informative value of single-point estimates.
@@
This uncertainty is due to the incomplete information provided by the finite-size HF and LF experimental designs, and can be reduced when more data is available. Therefore, it can be interpreted as epistemic uncertainty. Assuming that our regression model is able to accurately represent the underlying HF function in the presence of unlimited data, we have $\lim_{N_{H}\to\infty}(\psi_{\text{up},\alpha}(\boldsymbol{x}_{0})-\psi_{\text{lo},\alpha}(\boldsymbol{x}_{0}))=0$, meaning that as $N_{\text{H}}\to\infty$, the confidence interval for the regression model predictor collapses to a single value at each point.

### v2

# Uncertainty-aware multi-fidelity surrogate modeling with noisy data

@@ Abstract
Emulating high-accuracy computationally expensive models is crucial for tasks requiring numerous model evaluations, such as uncertainty quantification and optimization. When lower-fidelity models are available, they can be used to improve the predictions of high-fidelity models. Multi-fidelity surrogate models combine information from sources of varying fidelities to construct an efficient surrogate model. However, in real-world applications, uncertainty is present in both high- and low-fidelity models due to measurement or numerical noise, as well as lack of knowledge due to the limited experimental design budget. This paper introduces a comprehensive framework for multi-fidelity surrogate modeling that handles noise-contaminated data and is able to estimate the underlying noise-free high-fidelity model. Our methodology quantitatively incorporates the different types of uncertainty affecting the problem and emphasizes on delivering precise estimates of the uncertainty in its predictions both with respect to the underlying high-fidelity model and unseen noise-contaminated high-fidelity observations, presented through confidence and prediction intervals, respectively. Additionally, the proposed framework offers a natural approach to combining physical experiments and computational models by treating noisy experimental data as high-fidelity sources and white-box computational models as their low-fidelity counterparts. The effectiveness of our methodology is showcased through synthetic examples and a wind turbine application.
@@ 1 Introduction
Predicting the behavior of complex systems and quantifying the corresponding uncertainty is a ubiquitous challenge in engineering and applied sciences. To address this challenge, a variety of predictive models are employed. White-box computational models, typically implemented as computer simulations, use pre-existing knowledge on the underlying physics of a system to predict its behaviour. Conversely, black-box data-driven models act as global approximators of the response of a system, based on an available set of input-output observations (Rogers et al., 2017). In applications such as uncertainty quantification and optimization, which require numerous model evaluations, a specific subset of black-box models, known as surrogate models (SMs)—also known as metamodels or emulators—is often used to replace computationally expensive computational models. A SM acts as an inexpensive-to-evaluate approximator of an original model, and is constructed using a limited set of model evaluations, called the experimental design (ED), also known as training set in machine learning. Among the most widely used SMs for deterministic simulators are polynomial chaos expansions (PCE) (Xiu and Karniadakis, 2002; Blatman and Sudret, 2011), Gaussian processes (GPs) (Rasmussen and Williams, 2006), and support vector regression (Drucker et al., 1996).
@@
The choice of surrogate model is an integral part of the design and construction of a MFSM. Numerous MFSM techniques are based on Gaussian process modeling, following the autoregressive fusing scheme proposed by Kennedy and O’Hagan (2000). Such works include Forrester et al. (2007) and Le Gratiet and Garnier (2014), among others, with the latter reformulating the approach from Kennedy and O’Hagan (2000) to have a recursive form, allowing for a reduced computational complexity. Polynomial chaos expansion is another SM that has gained popularity in the past two decades for the purpose of MF surrogate modeling (Ng and Eldred, 2012; Palar et al., 2016). The approaches mentioned so far use linear information fusion, which entail assuming that a higher-fidelity response can be expressed as a linear combination of a lower-fidelity model and a discrepancy function. Recently, multi-fidelity modeling approaches have been proposed in the machine learning community, for example, the deep GP-based framework of Cutajar et al. (2018); Hebbal et al. (2021), the Bayesian neural network approaches from Meng et al. (2021); Kerleguer et al. (2024), and the generative adversarial network-based methodology proposed by Zhang et al. (2022). These approaches can capture the nonlinear relations between the different levels of fidelity. According to a comparison among different linear and nonlinear GP-based MF techniques performed by Brevault et al. (2020), when the high- and low-fidelity models are weakly correlated, nonlinear techniques can outperform linear and less complex techniques, with the caveat of requiring a larger quantity of HF data.
However, multi-fidelity methods do not necessarily involve the construction of a MFSM. Peherstorfer et al. (2018) classify MF methods into three categories: adaptation, where the LF model is enhanced with information from the HF model; fusion, which combines information from all fidelity levels simultaneously; and filtering, where the HF model is invoked only when the LF model is inaccurate or some criterion is met. Typically, multi-fidelity surrogate modeling methods fall under the adaptation (Kennedy and O’Hagan, 2000; Ng and Eldred, 2012; Chakraborty, 2021) or fusion (Bryson and Rumpfkeil, 2017; Yang et al., 2019) categories. In the context of uncertainty quantification, non-surrogate-based MF approaches include the control variate framework (Lavenberg and Welch, 1981; Gorodetsky et al., 2020; Pham and Gorodetsky, 2022), which falls under the fusion category, as well as importance sampling (Peherstorfer et al., 2016) and multistage Markov Chain Monte Carlo methods (Christen and Fox, 2005), both classified under filtering. Moreover, in some MF settings, a strict hierarchy of fidelity levels is not present. The MFNets framework proposed by Gorodetsky et al. (2020); Gorodetsky et al. (2021) provides a general approach to incorporate multiple fidelities with diverse relations.
@@
Most of the existing literature on MF surrogate modeling focuses on deterministic and noise-free high- and low-fidelity models. However, all real-world measurement devices have limited precision and resolution, and therefore, data resulting from measurements are generally contaminated by irreducible noise. In our context, this noise is considered as a source of aleatory uncertainty, as we assume that no more precise measurement device is available once a specific experimental data set has been provided. Moreover, the available training data for the construction of all the elements of a MFSM is in principle relatively small due to computational budget constraints. Thus, we consider all MFSM predictions as affected by epistemic uncertainty. Recent studies that consider the presence of noise in a MF setting include the work of Raissi et al. (2017), who use GP regression to infer the solutions of differential equations when noisy data of different fidelities are available. Furthermore, Zhang et al. (2018) demonstrate that their linear regression-based MF surrogate modeling technique is robust to noise in the HF data, and is also able to estimate the noise level when enough HF data is available. Also, Ficini et al. (2021) assess the robustness of a GP regression MFSM on problems affected by noisy objective function evaluations. Finally, few regression-based approaches have combined physical experiments and computer simulations using MFSMs (Kuya et al., 2011; Pandita et al., 2021).
@@
To the authors’ best knowledge, however, no work has yet introduced a comprehensive approach to multi-fidelity surrogate modeling that considers the presence of both noise and/or epistemic uncertainty in the high- and/or low-fidelity data. This approach should also provide a way to quantify the accuracy of the MF model predictions with respect to a) the underlying noise-free HF model and b) the noisy HF observations. The uncertainty about the mean prediction of a regression model is typically expressed via confidence intervals (CIs). The CIs express the epistemic uncertainty in the MFSM predictions. On the other hand, the uncertainty on the prediction of an unseen noise-contaminated observation is shown via prediction intervals (PIs) (Kutner et al., 2005). PIs consider the total of epistemic and aleatory uncertainty in the MFSM predictions, and thus, generally, they are wider than the corresponding CIs.
Many of the existing MF surrogate modeling methods do not provide uncertainty estimations about the MFSM predictions. When CIs are provided, their construction is usually linked to the particular methodology employed in constructing the MFSM, e.g., GP-based CIs (Raissi et al., 2017; Perdikaris et al., 2017). This can potentially make the frameworks less flexible when different MFSM architectures or SMs need to be explored. Additionally, GP-based confidence intervals often assume that the response distribution at each point is Gaussian, leading to symmetrical CIs around the MFSM response. However, this assumption does not always reflect reality. We propose a bootstrap-based approach for constructing CIs that does not assume any specific shape for the response distribution, making it more general and applicable to a wider range of scenarios. Moreover, no work has considered the distinction between confidence and prediction intervals in a multi-fidelity setting. Typically, prediction intervals are either disregarded or, mistakenly, confidence intervals are presented as prediction intervals.
@@
The goal of the present paper is to introduce a novel general framework for multi-fidelity surrogate modeling that is able to deal with noisy data and epistemic uncertainty due to limited training information. We assume that the noisy HF data originate from deterministic models contaminated by unbiased stochastic noise. Therefore, our MF framework aims at effectively emulating the underlying deterministic models, or in other words at denoising the noise-contaminated observations. An essential and original feature of our methodology is its ability to provide estimates of the different kinds of uncertainty in its predictions in the form of both confidence and prediction intervals.
Often, the response of a system can be predicted by one or more white-box computational models, while additional data can be obtained through physical experiments. The proposed multi-fidelity framework provides a direct approach for integrating both computational models and experimental data into a single surrogate model. Traditionally, physical experiments have been used to improve computational models through model calibration, where parameters of a model are inferred by fitting the model predictions to available experimental data (Kennedy and O'Hagan, 2001; Higdon et al., 2004; Pavlak et al., 2014). However, computational models may not fully capture the system’s complexity, making it important to incorporate data-driven elements. In this paper, we achieve this integration using multi-fidelity surrogate models. The computational models are assumed to capture the general physical behavior of the system, while the experiments can capture its entirety, but only on a very limited set of conditions, due to their associated costs. In particular, we consider the experimental data as noise-contaminated realizations of unknown HF models, whereas the available white-box computational models are considered as their LF counterparts. We approach white-box models within a non-intrusive context, eliminating the requirement for prior knowledge of the underlying equations. Other methods that have been explored in the literature to combine physical experiments and simulations include hybrid simulation, which combines physical and numerical substructures to create a hybrid model (Schellenberg et al., 2009; Abbiati et al., 2021), as well as scientific machine learning approaches, such as physics-informed neural networks (Raissi et al., 2019; Jagtap et al., 2022) and physics-informed PCE (Novák et al., 2024), which incorporate physical laws and constraints directly into the surrogate models.
@@
This paper is organized as follows: First, we recall the relevant theory for MFSMs and present our framework for MF surrogate modeling, including the special case of combining physical experiments and simulations. We then present the related confidence and prediction intervals, as well as the implementation details for each component of the framework. Subsequently, we assess the proposed method by applying it to two synthetic examples of increasing complexity, as well as a real-world application. Finally, we discuss concluding remarks and present prospects for future research.
@@ 2 Methods
In this section, we first formally state the multi-fidelity surrogate modeling problem, while establishing the notation we are adopting throughout the paper. Then, we describe our proposed methodology to combine experimental data and computational models in a multi-fidelity surrogate model. Subsequently, we propose to express the uncertainty related to the MFSM predictions using confidence and prediction intervals, and we discuss the interpretation of these intervals in the MF setting. Lastly, we outline our proposed implementation for constructing a MFSM, as well as for estimating its confidence and prediction intervals.
@@ 2.1 Multi-fidelity surrogate modeling
As implied in Equation 6, we assume a strict hierarchy of fidelities in our context, meaning that across the entire domain of definition, the higher-fidelity model is more accurate and/or more computationally expensive to evaluate than the lower-fidelity model. In addition, for the multi-fidelity surrogate model as in Equation 6 to outperform a surrogate model built solely from high-fidelity data, it is implicitly assumed that the discrepancy $\delta(\boldsymbol{x})$ between the HF model and the scaled LF model is less complex, and thus simpler to model, than the HF model itself.
Our MFSM approach to estimate the noise-free HF function $\psi_{\text{H}}(\boldsymbol{x})$ is based on regression; therefore, for the rest of this paper, the terms “multi-fidelity surrogate model" and “multi-fidelity regression model" are used interchangeably. The class of surrogates for the LF model and the discrepancy function $\delta(\boldsymbol{x})$ can be chosen among a wide range of regression models, including, among others, PCE, GP regression, or neural networks. However, the framework described in this section is independent of the particular choice of surrogate modeling methods, and hence, we will refrain from specifying the choice of SM for now.
@@
2.2 Combining physical experiments and computer simulations
Let us consider a set of observations $(\boldsymbol{{\mathcal{X}}}_{\text{H}},{\mathcal{Y}}_{\text{H}})$ obtained through an experimental campaign, each observation of which can be expressed as in Equation 3. Moreover, let us assume that the system response can also be predicted by a white-box computational model, such as a system of equations or a complex finite element simulator, denoted as ${\mathcal{M}}_{\text{L}}(\boldsymbol{x})$.
Then, we can assume that experiments can accurately capture the behavior of the system under investigation, but the information they provide is incomplete due to their scarcity. On the other hand, the available computational model can complement this information by providing a larger amount of data, but at the cost of lower accuracy, due either to inherent model simplifications (e.g. ignoring some physics), or to numerical limitations (e.g. discretization). To combine the two, we apply multi-fidelity surrogate modeling, wherein the experimental data is regarded as noise-affected realizations of a high-fidelity model, and the white-box computational model is treated as its low-fidelity equivalent. Following the rationale described in the previous section, the computational model is replaced by a surrogate model ^$\widehat{\mathcal{M}}_{\text{L}}(\boldsymbol{x})$, constructed with a set of model evaluations $(\boldsymbol{{\mathcal{X}}}_{\text{L}},{\mathcal{M}}_{\text{L}}(\boldsymbol{{\mathcal{X}}}_{\text{L}}))$.
Adopting the MF surrogate modeling information fusion scheme introduced previously, a multi-fidelity model predictor aiming to emulate the underlying noise-free HF function output $\psi_{\text{H}}(\boldsymbol{x})$ can be expressed as:
@@ 2.3 Confidence and prediction intervals
Confidence and prediction intervals are powerful means of conveying the uncertainty present in the predictions of a model, thus significantly increasing the informative value of single-point estimates.
@@
This uncertainty is due to the incomplete information provided by the finite-size HF and LF experimental designs, and can be reduced when more data is available. Therefore, it can be interpreted as epistemic uncertainty. Assuming that our multi-fidelity regression model is able to accurately represent the underlying HF function in the presence of unlimited data, we have $\lim_{N_{H}\to\infty}(\psi_{\text{up},\alpha}(\boldsymbol{x}_{0})-\psi_{\text{lo},\alpha}(\boldsymbol{x}_{0}))=0$, meaning that as $N_{\text{H}}\to\infty$, the confidence interval for the multi-fidelity regression model predictor collapses to a single value at each point.